OpenStudy (anonymous):

I doubt anyone knows, but I need to know step by step how to solve the trig equation: Secant [Sine-1 radical2/5] ....please help!

OpenStudy (amistre64):

the secant of sine inverse (sqrt(2)/5) is the question right? \(sec(sin(sqrt{2}/5)\)

OpenStudy (amistre64):

well: \(sin^{-1}(\sqrt{2}/5)\)

OpenStudy (amistre64):

you should be aware that the sin^-1 takes a number and gives back an angle between 0 and pi

OpenStudy (amistre64):

but easiest thing to do is make a triangle and draw a pic

OpenStudy (amistre64):

the answer is 5 according to the drawing

OpenStudy (anonymous):

my book says 5 radical 23 over 23.... i solve for my triangle taking sine(x)= radical 2 over 5 and find the third side. But from here, i dont know how to find secant

OpenStudy (amistre64):

secant = hyp/adj ; its just the inverse of cosine

OpenStudy (amistre64):

it is this right? \[sec(sin^{-1}(\frac{\sqrt{2}}{5})\]

OpenStudy (amistre64):

i did sqrt(2)^2 = 4 lol ....

OpenStudy (anonymous):

yes, i just dont understand how the answer is 5 radical 23 over 23

OpenStudy (amistre64):

sqrt(25 - 2) = sqrt(23) for the bottom part then right?

OpenStudy (anonymous):

yeah i dont understand how to get to that part. i just got my triangle and then i dont know where to go from there using secant

OpenStudy (amistre64):

if you know 2 sides of a right triangle you can find the thirs by the pythagoreum thrm

OpenStudy (amistre64):

(sqrt(2))^2 + base^2 = (5)^2 2 + base^2 = 25 base^2 = 25-2 base^2 = 23 base = sqrt(23)

OpenStudy (amistre64):

secant(A) = hyp/base sec = 5/sqrt(23)

OpenStudy (anonymous):

what is base?

OpenStudy (anonymous):

okay i got it, thanks!

OpenStudy (amistre64):

:) youre welcome :)

OpenStudy (anonymous):

I've posted another tricky one if you can solve it for me, that'd be greatly appreciated!